OpenStudy (anonymous):

Solve x^2-3x-10=0 by graphing.

OpenStudy (experimentx):

The following picture shows the graph if y = x^2-3x-10 http://www3.wolframalpha.com/Calculate/MSP/MSP22281a12h0c2biih523000005a6ebh997eg7d01c?MSPStoreType=image/gif&s=25&w=300&h=186&cdf=Coordinates&cdf=Tooltips choose points where line crosses y-axis

OpenStudy (experimentx):

the two red points...

OpenStudy (anonymous):

Thats when it's already graphed, how do you get to graphing it?

OpenStudy (anonymous):

x^2 -3x -10 = 0 (The first thing you want to do is find the roots.) => (x+2)(x-5) = 0 (so when y = 0, x = -2 or 5) => now you want the y -intercept, which is -10 in this case Finally, to get the minimum. differentiate the equation => dy/dx = 2x-3 = 0 so the minimum is at (3/2,-49/4) which is found by substituing x = 3/2 into the original equation. This allows you to sketch it as shown by experimentX

OpenStudy (experimentx):

for graphing put x=-8, calulate y then x=-7 calculate y, . ... .. go on until x=+8, and plot those points, you will have graph.

OpenStudy (anonymous):

Finding the derivative seems a bit beyond the level of this question. To find a maximum or minimum, using the equation -b/2a or completing the square makes more sense to me.